# Time-asymptotic self-similarity of the damped compressible Euler equations in parabolic scaling variables

### v1

@@ 3.3 Weak solutions satisfying (E2)
Let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ satisfy $\mathfrak{p}^{\prime}>0$. Let $(\rho,n)$ be a weak solution to (1.8) fulfilling (E2) with initial data $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}})^{2}$ such that ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\hat{\rho},0)<\infty$. Let $(\overline{\rho},\overline{n})\in C^{1}((0,\infty)\times{\mathbb{R}})^{2}\cap C([0,\infty)\times{\mathbb{R}})^{2}$ such that $\overline{\rho}>0$, or let $(\overline{\rho},\overline{n})\equiv(0,0)$ and additionally assume that $\mathfrak{p}$ satisfies (2.14). Further assume that $\overline{\rho}-\hat{\rho}$ and $\overline{n}$ have compact support in space, Then, the relative entropy inequality (3.4) holds.
@@
Moreover, there is $M>0$ such that $\overline{\rho}(\tau,y)-\rho_{\infty}(y)=\overline{n}(\tau,y)=0$ for $|y|>M$ and $\tau\in[0,T]$. Let $\psi\in C^{1}_{c}({\mathbb{R}})$ with $\psi=1$ in $[-M,M]$ and define $\chi(\tau,y)=\varphi(\tau)\psi(y)$. We use this test function in (3.15), (3.16) and (3.17), and we add the resulting identities and (3.22) to inequality (2.9), which yields
@@
Let $\phi\in C_{c}^{\infty}([0,\infty))$ be a smooth cut-off function with $\phi\equiv 1$ on $[0,1]$ and $\phi\equiv 0$ on $[2,\infty)$, and set $\phi_{R}(y):=\phi(|y|/R)$ for $R>0$. We approximate the reference pair $(\overline{\rho},\overline{n})$ by
@@
$\mathrm{e}^{-\tau}\frac{\rho}{2}\Big{|}\frac{n}{\rho}-\frac{\overline{n}_{R}}{\overline{\rho}_{R}}\Big{|}^{2}=\mathrm{e}^{-\tau}\frac{\rho}{2}\Big{|}\frac{n}{\rho}-\frac{\overline{n}}{\overline{\rho}}\Big{|}^{2}-\mathrm{e}^{-\tau}(\phi_{R}-1)n\frac{\overline{n}}{\overline{\rho}}+\mathrm{e}^{-\tau}(\phi_{R}^{2}-1)\rho\frac{\overline{n}^{2}}{\overline{\rho}^{2}},$
@@
$-\int_{0}^{1}h^{\prime\prime}\big{(}\overline{\rho}+\theta(1-\phi_{2R})(\rho_{\infty}-\overline{\rho})\big{)}\big{(}\rho-\overline{\rho}-\theta(1-\phi_{2R})(\rho_{\infty}-\overline{\rho})\big{)}(1-\phi_{2R})(\rho_{\infty}-\overline{\rho})\,\mathrm{d}\theta$
@@
${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\overline{\rho}_{R}(0),\overline{n}_{R}(0))$ $\to{\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\overline{\rho}(0),\overline{n}(0))$ as $\quad\text{as }R\to\infty,$
${\mathcal{E}}(\tau;\rho,n\,|\,\overline{\rho}_{R},\overline{n}_{R})$ $\to{\mathcal{E}}(0;\rho,n\,|\,\overline{\rho},\overline{n})$ in $\text{in }L^{\infty}_{\mathrm{loc}}([0,\infty))$ as $\quad\text{as }R\to\infty.$
@@
Let us compare the two main results on the relative entropy inequality from Theorem 3.11 and Theorem 3.14 under the different criteria (E1) and (E2). While (E1) is slightly more restrictive than (E2), see Remark 3.10, the corresponding assumption (3.21) on the pair $(\overline{\rho},\overline{n})$ is slightly more general than (3.24). Although both assumptions (3.21) and (3.24) are clearly not optimal, they are sufficient for our purposes.
@@ 3.4 Weak-strong uniqueness
$\int_{0}^{\tau+\delta}-{\omega_{\delta}^{\prime}}(\sigma)\,\mathrm{e}^{-\int_{0}^{\sigma}a(\nu)\,\mathrm{d}\nu}E(\sigma)\,\mathrm{d}\sigma\leq E_{0}+\int_{0}^{\tau+\delta}\omega_{\delta}(\sigma)\,\mathrm{e}^{-\int_{0}^{\sigma}a(\nu)\,\mathrm{d}\nu}b(\sigma)\,\mathrm{d}\sigma.$
@@
$\mathrm{e}^{-\int_{0}^{\tau}a(\nu)\,\mathrm{d}\nu}E(\tau)\leq E_{0}+\int_{0}^{\tau}b(\sigma)\,\mathrm{e}^{-\int_{0}^{\sigma}a(\nu)\,\mathrm{d}\nu}\,\mathrm{d}\sigma,$
@@ 4 Convergence towards similarity profiles
In this section we prove our main results on the convergence of weak solutions $(\rho,n)$ to the scaled system (1.8) towards the similarity profile $(\rho^{*},n^{*})$ defined by (1.11). To this end, we consider the residuals $\overline{R}_{i}$ as defined in (3.1), and the function $\xi$ given in (3.2), all with respect to the profile $(\rho^{*},n^{*})$ and denote the residuals with $R_{i}^{*}$.
@@ 4.1 Coincident limits at spatial infinity
We first consider the case when the limit densities $\rho_{\pm}$ coincide, so that $\hat{\rho}=\rho_{+}=\rho_{-}\geq 0$. Then, the corresponding similarity profile is constant and given by $(\rho^{*},n^{*})(y)\equiv(\hat{\rho},0)$. In particular, we have $R^{*}_{1}=R^{*}_{2}=0$ and $\xi=0$. Utilizing inequality (3.4) for the total relative entropy, we deduce the convergence result in the case of coincident limits.
@@
for a.a. $\tau,\sigma\in(0,\infty)$ with $\tau<\sigma$. Omitting the entropy terms on the left-hand side, and using the decay estimate from Theorem 2.4, we obtain (2.16). ∎
@@ 4.2 Different limits at spatial infinity
Now we address the case where the limit densities are both positive and do not coincide, that is, $\rho_{\pm}>0$ and $\rho_{+}\neq\rho_{-}$. Hence, the unique similarity profile $(\rho^{*},n^{*})$ solving (1.11) is not constant. Further, by using that $\rho^{*}$ and $n^{*}$ are time-independent, the corresponding residuals $R_{i}^{*}$ are given by $R^{*}_{1}=0$ and $R^{*}_{2}=R^{*}$ with $R^{*}$ defined in (2.12). This in turn leads to simplified expressions for $\xi_{j}$ from (3.2) as we now have
@@
$\leq\mathrm{e}^{-(\frac{1}{2}{-}\theta)\tau+\frac{1}{2}\mu}\Big{(}{\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\rho^{*},n^{*})+\frac{K}{\theta}\Big{)}+\frac{K}{2}\mathrm{e}^{-\tau/2}$
@@ 5 Extension to multiple dimensions
$\widetilde{\rho}_{t}+\operatorname{div}\widetilde{m}$ $=0,$ (5.1)
$\widetilde{m}_{t}+\operatorname{div}\Big{(}\frac{\widetilde{m}\otimes\widetilde{m}}{\widetilde{\rho}}+\mathfrak{p}(\widetilde{\rho}){\mathbb{I}}\Big{)}+\alpha\widetilde{m}$ $=0$
@@
Instead of two limit densities $\rho_{\pm}$, we now prescribe a positive angular function
@@ 5.1 The notion of energy-variational solutions
Let $(\widetilde{\rho}_{0},\widetilde{m}_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d};{\mathbb{R}}^{d})$ such that ${\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$. We call $(\widetilde{\rho},\widetilde{m},\widetilde{E})\in L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d};{\mathbb{R}}^{d})\times\mathrm{BV}([0,\infty))$ an energy-variational solution to (5.1) if there is $\widetilde{\rho}_{\infty}\in C^{1}([0,\infty)\times{\mathbb{R}}^{d})$ with (5.3) such that ~$\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)\leq\widetilde{E}$ a.e. in $(0,\infty)$, and if there exists a $1$-homogeneous regularity weight $\widetilde{\mathcal{K}}\colon C_{c}^{1}({\mathbb{R}}^{d};{\mathbb{R}}^{d})\to[0,\infty)$ such that
@@
As before, with the assumption ${\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$ we implicitly assume $\rho_{0}\geq 0$ a.e. Furthermore, the statement and proof of Proposition 3.3 can immediately be transferred to the multi-dimensional setting. Therefore, ${\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$ is equivalent to ${\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\widetilde{\rho}_{\infty}(0),0)<\infty$.
@@
for all $\phi,\psi$ and $s<t$ as above. This can be seen by a classical variational argument: Choose test functions $\mu\phi$ and $\lambda\psi$ for $\mu,\lambda\geq 0$ and consider the limit cases $\mu=0$, $\lambda=0$, $\mu\to\infty$ and $\lambda\to\infty$. While (5.5a) corresponds to a standard weak formulation of the continuity equation, the main difference between weak solution and energy-variational solutions to (5.1) is visible in (5.5b) and (5.5c). The energy variable $\widetilde{E}$ occurs in both equations as an additional auxiliary variable. The energy defect $\widetilde{E}-\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)\geq 0$ between $\widetilde{E}$ and the mechanical energy ~$\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)$ accounts for the occurrence of anomalous dissipation due to turbulent effects. Inequality (5.5b) relaxes the momentum equation to an inequality, where the right-hand side measures the deviation from a weak form in terms of this energy defect multiplied by the regularity weight $\widetilde{\mathcal{K}}$. Moreover, (5.5c) corresponds to the (relative) entropy inequality, where the mechanical energy is replaced with $\widetilde{E}$. In particular, if $\widetilde{E}=\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)$, then (5.5) reduces to a classical weak formulation of (1.1) combined with an energy inequality.
@@
In order to obtain a finite (relative) energy ${\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)$, the existence of a specific reference density $\widetilde{\rho}_{\infty}$ with suitable values as $|x|\to\infty$ is necessary. However, the state variable $(\widetilde{\rho},\widetilde{m})$ of an energy-variational solution $(\widetilde{\rho},\widetilde{m},\widetilde{E})$ is independent of this choice. Indeed, if we have two reference densities $\rho^{1}_{\infty}$ and $\rho^{2}_{\infty}$, then $(\widetilde{\rho},\widetilde{m},E^{1})$ is an energy-variational solution to (1.1) for $\widetilde{\rho}_{\infty}=\rho^{1}_{\infty}$ if and only if $(\widetilde{\rho},\widetilde{m},E^{2})$ is an energy-variational solution for $\widetilde{\rho}_{\infty}=\rho_{\infty}^{2}$, where
@@
~$\widetilde{\mathcal{E}}(\rho,m\,|\,\rho^{2}_{\infty},0)=\widetilde{\mathcal{E}}(\rho,m\,|\,\rho^{1}_{\infty},0)+\int_{{\mathbb{R}}^{d}}h(\rho^{1}_{\infty}\,|\,\rho^{2}_{\infty})+\big{(}h^{\prime}(\rho^{1}_{\infty})-h^{\prime}(\rho^{2}_{\infty})\big{)}(\widetilde{\rho}-\rho^{1}_{\infty})\,\mathrm{d}x,$
and by substituting the test function $\phi$ with $\phi+h^{\prime}(\rho^{1}_{\infty})-h^{\prime}(\rho^{2}_{\infty})$, which is admissible since, as before, it suffices to consider Lipschitz continuous test functions.
For the sake of completeness, let us show that, in contrast to weak solutions, the global-in-time existence of energy-variational solutions to (5.1) can be shown in the multi-dimensional framework. To this end, we use the general existence result from [ALR24], which allows for the presence of dissipative terms.
@@
Let $\alpha\geq 0$ and $\mathfrak{p}(\rho)=k\rho^{\gamma}$ for $k>0$, $\gamma>1$. Let $\hat{\rho}$ as in (5.2) such that either $\hat{\rho}>0$ or $\hat{\rho}\equiv 0$. Let $(\widetilde{\rho}_{0},\widetilde{m}_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d};{\mathbb{R}}^{d})$ such that ~$\widetilde{\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$. Then there exists an energy-variational solution $(\widetilde{\rho},\widetilde{m},\widetilde{E})$ to (5.1) with
@@
with $\widetilde{E}(0)={\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)$. The corresponding regularity weight $\widetilde{\mathcal{K}}$ is given by (5.6).
@@
The dissipation functional is given by
@@
where we set $\frac{|\widetilde{m}|^{2}}{\rho}=+\infty$ if $\rho<0$, and if $\rho=0$ and $m\neq 0$. We choose the regularity weight $\widetilde{\mathcal{K}}$ as in (5.6), and show that all assumptions in the abstract existence result from [ALR24, Theorem 3.3] are satisfied. In particular, since $\eta$ is strictly convex and lower semicontinuous, this also applies to the associated relative entropy, so that ${\mathcal{F}}$ is strictly convex and lower semicontinuous. To show coercivity of ${\mathcal{F}}$, let $K=\widetilde{\rho}_{\infty}({\mathbb{R}}^{d})$. Then Lemma 3.4 yields a constant $r_{0}\geq 0$ such that (3.11) holds for all $\overline{\rho}=\widetilde{\rho}_{\infty}(x)$, $x\in{\mathbb{R}}^{d}$. For $(\widetilde{\sigma},\widetilde{m})\in{\mathbb{V}}$, we introduce the sets

### v2

@@ 3.3 Weak solutions satisfying (E2)
Let $\mathfrak{p}\in C([0,\infty))\cap C^{1}(0,\infty)$ satisfy $\mathfrak{p}^{\prime}>0$. Let $(\rho,n)$ be a weak solution to (1.8) fulfilling (E2) with initial data $(\rho_{0},n_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}})^{2}$ such that ${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\hat{\rho},0)<\infty$. Let either $(\overline{\rho},\overline{n})\in C^{1}((0,\infty)\times{\mathbb{R}})^{2}\cap C([0,\infty)\times{\mathbb{R}})^{2}$ such that $\overline{\rho}>0$, or let $(\overline{\rho},\overline{n})\equiv(0,0)$ and assume that $\mathfrak{p}$ satisfies (2.14). Further assume that $\overline{\rho}-\hat{\rho}$ and $\overline{n}$ have compact support in space. Then the relative entropy inequality (3.4) holds.
@@
[…]
@@
Let $\phi\in C_{c}^{\infty}([0,\infty))$ be a smooth cut-off function with $\phi\equiv 1$ on [0,1][0,1][ 0 , 1 ] and $\phi\equiv 0$ on [2,)[2$\infty$)[ 2 , ∞ ), and set $\phi_{R}(y):=\phi(|y|/R)$ for $R>0$. We approximate the reference pair $(\overline{\rho},\overline{n})$ by
@@
$\mathrm{e}^{-\tau}\frac{\rho}{2}\Big{|}\frac{n}{\rho}-\frac{\overline{n}_{R}}{\overline{\rho}_{R}}\Big{|}^{2}=\mathrm{e}^{-\tau}\frac{\rho}{2}\Big{|}\frac{n}{\rho}-\frac{\overline{n}}{\overline{\rho}}\Big{|}^{2}-\mathrm{e}^{-\tau}(\phi_{R}-1)n\frac{\overline{n}}{\overline{\rho}}+\mathrm{e}^{-\tau}(\phi_{R}^{2}-1)\frac{\rho}{2}\frac{\overline{n}^{2}}{\overline{\rho}^{2}},$
@@
$-\int_{0}^{1}h^{\prime\prime}\big{(}\overline{\rho}+\theta(1-\phi_{2R})(\rho_{\infty}-\overline{\rho})\big{)}\big{(}\rho-\overline{\rho}-\theta(1-\phi_{2R})(\rho_{\infty}-\overline{\rho})\big{)}(1-\phi_{2R})(\rho_{\infty}-\overline{\rho})\,\mathrm{d}\theta,$
@@
${\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\overline{\rho}_{R}(0),\overline{n}_{R}(0))$ $\to{\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\overline{\rho}(0),\overline{n}(0))$ $\quad\text{as }R\to\infty,$
${\mathcal{E}}(\tau;\rho,n\,|\,\overline{\rho}_{R},\overline{n}_{R})$ $\to{\mathcal{E}}(0;\rho,n\,|\,\overline{\rho},\overline{n})$ $\text{in }L^{\infty}_{\mathrm{loc}}([0,\infty))$ $\quad\text{as }R\to\infty.$
@@
Let us compare the two main results on the relative entropy inequality from Theorem 3.11 and Theorem 3.14 under the different criteria (E1) and (E2). While (E1) is more restrictive than (E2), see Remark 3.10, the corresponding assumption (3.21) on the pair $(\overline{\rho},\overline{n})$ is slightly more general than (3.24). Although both assumptions (3.21) and (3.24) are clearly not optimal, they are sufficient for our purposes.
@@ 3.4 Weak-strong uniqueness
$\int_{0}^{\tau+\delta}-{\omega_{\delta}^{\prime}}(\sigma)\,\mathrm{e}^{-\int_{0}^{\sigma}a(\nu)\,\mathrm{d}\nu}E(\sigma)\,\mathrm{d}\sigma\leq E_{0}+\int_{0}^{\tau+\delta}\omega_{\delta}(\sigma)\,b(\sigma)\,\mathrm{e}^{-\int_{0}^{\sigma}a(\nu)\,\mathrm{d}\nu}\,\mathrm{d}\sigma.$
@@
$\mathrm{e}^{-\int_{0}^{\tau}a(\nu)\,\mathrm{d}\nu}E(\tau)\leq E_{0}+\int_{0}^{\tau}b(\sigma)\,\mathrm{e}^{-\int_{0}^{\sigma}a(\nu)\,\mathrm{d}\nu}\,\mathrm{d}\sigma\quad\text{for a.e.~}\tau>0,$
@@ 4 Convergence towards similarity profiles
In this section we prove our main results in the one-dimensional case, namely the convergence of weak solutions $(\rho,n)$ to the scaled system (1.8) towards the similarity profile $(\rho^{*},n^{*})$ defined by (1.11). To this end, we consider the residuals $\overline{R}_{i}$ as defined in (3.1), and the function $\xi$ given in (3.2), all with respect to the profile $(\rho^{*},n^{*})$ and denote the residuals with $R_{i}^{*}$.
@@ 4.1 Coincident limits at spatial infinity
We first consider the case when the limits $\rho_{\pm}$ coincide, so that $\hat{\rho}=\rho_{+}=\rho_{-}\geq 0$. Then, the corresponding similarity profile is constant and given by $(\rho^{*},n^{*})(y)\equiv(\hat{\rho},0)$. In particular, we have $R^{*}_{1}=R^{*}_{2}=0$ and $\xi=0$. Utilizing inequality (3.4) for the total relative entropy, we deduce the convergence result in the case of coincident limits.
@@
for a.a. $\tau,\sigma\in(0,\infty)$ with $\tau<\sigma$. Omitting the entropy terms on the left-hand side, and using the decay estimate (2.15), we obtain (2.16). ∎
@@ 4.2 Different limits at spatial infinity
Now we address the case where the limits $\rho_{\pm}$ are both positive and do not coincide, that is, $\rho_{\pm}>0$ and $\rho_{+}\neq\rho_{-}$. Hence, the unique similarity profile $(\rho^{*},n^{*})$ solving (1.11) is not constant. Further, by using that $\rho^{*}$ and $n^{*}$ are time-independent, the corresponding residuals $R_{i}^{*}$ are given by $R^{*}_{1}=0$ and $R^{*}_{2}=R^{*}$ with $R^{*}$ defined in (2.12). This in turn leads to simplified expressions for $\xi_{j}$ from (3.2) as we now have
@@
$\leq\mathrm{e}^{-(\frac{1}{2}{-}\theta)\tau+\frac{1}{2}\mu}\Big{(}{\mathcal{E}}(0;\rho_{0},n_{0}\,|\,\rho^{*},n^{*})+\frac{K}{\theta}\Big{)}+2K\mathrm{e}^{-\tau/2}$
@@ 5 Extension to multiple dimensions
$\widetilde{\rho}_{t}+\operatorname{div}\widetilde{m}$ =0=0,= 0 , (5.1)
$\widetilde{m}_{t}+\operatorname{div}\Big{(}\frac{\widetilde{m}\otimes\widetilde{m}}{\widetilde{\rho}}+\mathfrak{p}(\widetilde{\rho}){\mathbb{I}}\Big{)}+\alpha\widetilde{m}$ =0=0= 0
@@
Instead of two different limits $\rho_{\pm}$, we now prescribe an angular function
@@ 5.1 The notion of energy-variational solutions
Let $(\widetilde{\rho}_{0},\widetilde{m}_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d};{\mathbb{R}}^{d})$ such that ~$\widetilde{\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$. We call $(\widetilde{\rho},\widetilde{m},\widetilde{E})\in L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}([0,\infty)\times{\mathbb{R}}^{d};{\mathbb{R}}^{d})\times\mathrm{BV}([0,\infty))$ an energy-variational solution to (5.1) if there is $\widetilde{\rho}_{\infty}\in C^{1}([0,\infty)\times{\mathbb{R}}^{d})$ with (5.3) such that ~$\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)\leq\widetilde{E}$ a.e. in (0,)(0$\infty$)( 0 , ∞ ), and if there exists a 111-homogeneous regularity weight $\widetilde{\mathcal{K}}\colon C_{c}^{1}({\mathbb{R}}^{d};{\mathbb{R}}^{d})\to[0,\infty)$ such that
@@
As before, with the assumption ~$\widetilde{\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$ we implicitly assume $\rho_{0}\geq 0$ a.e. Furthermore, the statement and proof of Proposition 3.3 can immediately be transferred to the multi-dimensional setting. Therefore, ~$\widetilde{\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$ is equivalent to ~$\widetilde{\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\widetilde{\rho}_{\infty}(0),0)<\infty$.
@@
for all $\phi,\psi$ and $s<t$ as above. This can be seen by a classical variational argument: Choose test functions $\mu\phi$ and $\lambda\psi$ for $\mu,\lambda\geq 0$ and consider the limit cases $\mu=0$, $\lambda=0$, $\mu\to\infty$ and $\lambda\to\infty$. While (5.5a) corresponds to a standard weak formulation of the continuity equation, the main difference between weak solutions and energy-variational solutions to (5.1) is visible in (5.5b) and (5.5c). The energy variable $\widetilde{E}$ occurs in both equations as an additional auxiliary variable. Inequality (5.5b) relaxes the momentum equation to an inequality, where the right-hand side measures the deviation from a weak form in terms of the energy defect $\widetilde{E}-\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)\geq 0$ multiplied by the regularity weight $\widetilde{\mathcal{K}}$. Moreover, (5.5c) corresponds to the (relative) entropy inequality, where the mechanical energy is replaced with $\widetilde{E}$. In particular, if $\widetilde{E}=\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)$, then (5.5) reduces to a classical weak formulation of (5.1) combined with an energy inequality.
@@
In order to obtain a finite (relative) energy ~$\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\widetilde{\rho}_{\infty},0)$, the existence of a specific reference density $\widetilde{\rho}_{\infty}$ with suitable values as $|x|\to\infty$ is necessary. However, the state variable $(\widetilde{\rho},\widetilde{m})$ of an energy-variational solution $(\widetilde{\rho},\widetilde{m},\widetilde{E})$ is independent of this choice. Indeed, if we have two reference densities $\rho^{1}_{\infty}$ and $\rho^{2}_{\infty}$, then $(\widetilde{\rho},\widetilde{m},E^{1})$ is an energy-variational solution to (5.1) for $\widetilde{\rho}_{\infty}=\rho^{1}_{\infty}$ if and only if $(\widetilde{\rho},\widetilde{m},E^{2})$ is an energy-variational solution for $\widetilde{\rho}_{\infty}=\rho_{\infty}^{2}$, where
@@
~$\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\rho^{2}_{\infty},0)=\widetilde{\mathcal{E}}(\widetilde{\rho},\widetilde{m}\,|\,\rho^{1}_{\infty},0)+\int_{{\mathbb{R}}^{d}}h(\rho^{1}_{\infty}\,|\,\rho^{2}_{\infty})+\big{(}h^{\prime}(\rho^{1}_{\infty})-h^{\prime}(\rho^{2}_{\infty})\big{)}(\widetilde{\rho}-\rho^{1}_{\infty})\,\mathrm{d}x,$
and by substituting the test function $\phi$ with $\phi+h^{\prime}(\rho^{1}_{\infty})-h^{\prime}(\rho^{2}_{\infty})$.
For the sake of completeness, let us show that, in contrast to weak solutions, the global-in-time existence of energy-variational solutions to (5.1) can be shown in the multi-dimensional framework. To this end, we use the general existence result from [ALR24], which allows for the presence of dissipative terms. Here we denote the class of weakly continuous functions from [0,)[0$\infty$)[ 0 , ∞ ) to a Banach space $X$ by $C_{w}([0,\infty);X)$.
@@
Let $\alpha\geq 0$ and $\mathfrak{p}(\rho)=k\rho^{\gamma}$ for $k>0$, $\gamma>1$. Let either $\hat{\rho}$ be as in (5.2) or $\hat{\rho}\equiv 0$. Let $(\widetilde{\rho}_{0},\widetilde{m}_{0})\in L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d})\times L^{1}_{\mathrm{loc}}({\mathbb{R}}^{d};{\mathbb{R}}^{d})$ such that ~$\widetilde{\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)<\infty$. Then there exists an energy-variational solution $(\widetilde{\rho},\widetilde{m},\widetilde{E})$ to (5.1) with
@@
with $\widetilde{E}(0)=\widetilde{\mathcal{E}}(\widetilde{\rho}_{0},\widetilde{m}_{0}\,|\,\hat{\rho},0)$. The corresponding regularity weight $\widetilde{\mathcal{K}}$ is given by (5.6).
@@
Since $\widetilde{\rho}_{\infty}$ is time-independent, the dissipation functional is given by
@@
where we set $\frac{|m|^{2}}{\rho}=+\infty$ if $\rho<0$, and if $\rho=0$ and $m\neq 0$. We choose the regularity weight $\widetilde{\mathcal{K}}$ as in (5.6), and show that all assumptions in the abstract existence result from [ALR24, Theorem 3.3] are satisfied. In particular, since $\eta$ is strictly convex and lower semicontinuous, this also applies to the associated relative entropy, so that ${\mathcal{F}}$ is strictly convex and lower semicontinuous. To show coercivity of ${\mathcal{F}}$, let $K=\widetilde{\rho}_{\infty}({\mathbb{R}}^{d})$. Then Lemma 3.4 yields a constant $r_{0}\geq 0$ such that (3.11) holds for all $\overline{\rho}=\widetilde{\rho}_{\infty}(x)$, $x\in{\mathbb{R}}^{d}$. For $(\widetilde{\sigma},\widetilde{m})\in{\mathbb{V}}$, we introduce the sets